OpenStudy (anonymous):

Which phrase correctly orders –34, 15, –23, and 2? A. 2 < 15 < –34 < –23 B. –34 < –23 < 15 < 2 C. 2 < 15 < –23 < –34 D. –34 < –23 < 2 < 15

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

youre right

OpenStudy (anonymous):

cool&thx

OpenStudy (anonymous):

i'll tell u me score

OpenStudy (anonymous):

wait

OpenStudy (cresciez):

It's B lol

OpenStudy (anonymous):

its d i forgot negatives go before the posit

OpenStudy (anonymous):

positives

OpenStudy (cresciez):

Negatives come first - despite their big numbers, the negative sign is a big hint. The bigger the negative number, the lesser it is.

OpenStudy (anonymous):

well i got a 76%

OpenStudy (cresciez):

Oh yeah, it is D oops haha, I just paid attention to the negatives

OpenStudy (cresciez):

Still a lovely score ^.^

OpenStudy (anonymous):

lol yea i was only looking at the numbers im really sorry @Crazy_questions

OpenStudy (anonymous):

yea but i need at lest a 80% to pass

OpenStudy (cresciez):

Can you retake the test rn?

OpenStudy (anonymous):

yea can you?

OpenStudy (anonymous):

yea ican

OpenStudy (anonymous):

ok retry then this time your gonna get your 80 or up

OpenStudy (anonymous):

ok :)

OpenStudy (cresciez):

Awesome - I'll be happy to help you with the problems if you need any. :^)

OpenStudy (anonymous):

ok question 1 –32 + 22 A. –10 B. –54 C. 54 D. 10 lol so easy it -10

OpenStudy (cresciez):

You're correct~

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok question 3 b + 6.7 = 19.4 there is so post it be a box here

OpenStudy (anonymous):

question 2 was easy

OpenStudy (anonymous):

12.7

OpenStudy (cresciez):

^ Mhm :>

OpenStudy (anonymous):

ok question 4 Nancy drove from New York to Boston, a distance of 1144.97 km. She drove with a friend for the first 711.5 km, and then by herself for the remaining distance. How far did Nancy drive by herself? A. 1856.47 km B. 1855.47 km C. 333.47 km D. 433.47 km

OpenStudy (anonymous):

i got d

OpenStudy (anonymous):

ok question 5 Marcus biked 15.7 miles on Saturday and 13.9 miles the following Thursday. Ellen biked the same total distance as Marcus, but only biked 12.8 miles on Saturday. How far did Ellen bike on Thursday? A. 29.6 miles B. 16.8 miles C. 2.9 miles D. 28.5 miles

OpenStudy (anonymous):

b

OpenStudy (cresciez):

I think B - she had to make up on the leftover miles

OpenStudy (anonymous):

we all pick b ok q 6 lol esay 1 again What is the opposite of 93? A. 93 B. –39 C. –93 D. 39 its c

OpenStudy (cresciez):

^ C

OpenStudy (cresciez):

Correct again :>

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ok q7 Which phrases correctly describe the location of –52 on the number line? Choose exactly two answers that are correct. A. to the left of –38 B. to the right of –51 C. to the right of –63 D. to the left of –76

OpenStudy (cresciez):

B and D

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

ok q8 How far is point P from point Q? Number line -10 to +10 - assessment graphic A. 8 units B. 9 units C. 10 units D. 7 units

OpenStudy (cresciez):

7 units, I believe; so D ^-^

OpenStudy (cresciez):

No wait A

OpenStudy (anonymous):

i say a

OpenStudy (cresciez):

I didn't realize that you should count each point

OpenStudy (anonymous):

ok q9 What is the perimeter of this triangle? (box go here) m Image of triangle with sides labelled with 6 m, 3.2 m, and 7.6 m

OpenStudy (cresciez):

Perimeter is 16.8; just add the numbers w/ a calculator to get the total perimeter

OpenStudy (anonymous):

nice thats correct

OpenStudy (cresciez):

:3

OpenStudy (anonymous):

ok q 10 Anne walked 3.2 kilometers. How many meters did Anne walk? A. 320 meters B. 0.0032 meters C. 3200 meters D. 0.032 meters

OpenStudy (cresciez):

13,200 I think - C

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ok q 11 Which unit would be most appropriate for measuring the width of a house? A. mm B. km C. cm D. m

OpenStudy (anonymous):

d

OpenStudy (cresciez):

^ feet or meters is the best unit to use

OpenStudy (anonymous):

ok q12 Which equation could be used to find the perimeter of an isosceles triangle, with two sides 5 inches long, a base 6 inches long and a height of 4 inches? A. P = (6 × 4) ÷ 2 B. P = 2 (5 + 6) C. P = 5 + 6 + 4 D. P = 2 × 5 + 6

OpenStudy (cresciez):

I would say C - totaling perimeter is just adding

OpenStudy (anonymous):

correct

OpenStudy (cresciez):

But I could be wrong; because it says there are TWO sides 5 inches long

OpenStudy (cresciez):

Let's think about it for a sec

OpenStudy (anonymous):

q13 Which phrase correctly orders –23, 11, –17, and 5? A. 5 < 11 < –17 < –23 B. –23 < –17 < 5 < 11 C. –23 < 11 < –17 < 5 D. –17 < –23 < 5 < 11 A??

OpenStudy (cresciez):

It's D on q. 12 - I researched it ^.^

OpenStudy (cresciez):

D for sure - so please change that one

OpenStudy (anonymous):

ok

OpenStudy (cresciez):

B is for q 13

OpenStudy (cresciez):

'cause see - the negatives are the smallest; they're lower than 0

OpenStudy (anonymous):

b

OpenStudy (cresciez):

the bigger the number of a negative, the farther away from 0, so that automatically comes before any positive number

OpenStudy (anonymous):

ok i'll tell u my score

OpenStudy (anonymous):

it should be gud

OpenStudy (anonymous):

:(

OpenStudy (cresciez):

YOU'RE KIDDING

OpenStudy (anonymous):

JK

OpenStudy (cresciez):

if you got below 80 that's a screwy test

OpenStudy (cresciez):

facepalm

OpenStudy (anonymous):

omg ur kidding me like fr

OpenStudy (anonymous):

i got a 84%

OpenStudy (cresciez):

YAYYYYY

OpenStudy (cresciez):

Imma help the others ^_^

OpenStudy (anonymous):

lol

OpenStudy (cresciez):

Byeee

OpenStudy (anonymous):

lol niceeeeee 

OpenStudy (anonymous):

i press 9 key i was so post to press 0 key for this face :)

OpenStudy (anonymous):

O LOL well wen u need help jus tag me

OpenStudy (cresciez):

Same :>

OpenStudy (anonymous):

thank u so much 4 hlping

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

:)